# CLoMo: Counterfactual Logical Modification with Large Language Models

### v3

@@ Abstract
In this study, we delve into the realm of counterfactual reasoning capabilities of large language models (LLMs). Our primary objective is to cultivate the counterfactual thought processes within LLMs and rigorously assess these processes for their validity. Specifically, we introduce a novel task, Counterfactual Logical Modification (CLoMo), and a high-quality human-annotated benchmark. In this task, LLMs must adeptly alter a given argumentative text to uphold a predetermined logical relationship. To effectively evaluate a generation model’s counterfactual capabilities, we propose an innovative evaluation metric, the decomposed Self-Evaluation Score (SES) to directly evaluate the natural language output of LLMs instead of modeling the task as a multiple-choice problem. Analysis shows that the proposed automatic metric aligns well with human preference. Our experimental results show that while LLMs demonstrate a notable capacity for logical counterfactual thinking, there remains a discernible gap between their current abilities and human performance. We will release the data and code after publication.
@@ 1 Introduction
In our study, we delve into the realm of evaluating large language models’ (LLMs) ability to generate counterfactually coherent thoughts. Specifically, we proposed an innovative evaluation system that quantitatively measures the evolution of information in statement pairs, ensuring that they adhere to a specified logical relationship. Our approach includes designing a specialized task where models are presented with mismatched argument-premise pairs bound by a specific logical relation. The objective for these models is to adeptly modify the argument text until the specified logical relation is satisfactorily established. In conjunction with this task, we have created the first dataset of its kind, comprising dual argument-premise pairs, each annotated with a defined logical relation. This dataset is vital for facilitating logically restricted counterfactual modifications, and we have enriched it with human-written modifications to serve as a benchmark for evaluation.
Our experimental investigations encompass a range of large language models, including GPT-4 OpenAI (2023) and GPT-3.5-Turbo OpenAI (2022), as well as smaller models from the LLaMA Touvron et al. (2023a) and LLaMA 2 Touvron et al. (2023b) families. Through these experiments, we have discerned that the task of CLoMo poses a significant challenge. It becomes evident that these models’ current counterfactual logical reasoning capabilities fall short of the desired proficiency. This observation underscores the need for further advancements in enhancing the counterfactual reasoning abilities of existing language models, paving the way for more sophisticated and logically coherent AI systems.
@@
  • We propose the task of Counterfactual Logical Modification and contribute a corresponding CLoMo. to evaluate the counterfactual reasoning capability of LLMs in the scenario of complicated textual logical reasoning.
@@
  • We conduct experiments on representative LLMs (GPT-3.5, GPT-4) and small language models (the LLaMA and LLaMA 2 families) and find that CLoMo is a very challenging task and the counterfactual logical reasoning ability of the existing model needs to be improved.
@@
Necessary Assumption (38.5%)
Argument: Statement1: Journalist: the advice of social scientists is frequently overlooked by politicians making social policy. Statement 2: Because it is not unreasonable to discount scientific assertions backed by weak evidence, politicians should not generally be criticized for ignoring social science, for social scientists, unlike physical scientists, seldom agree on the claims made even within their own specialty.
@@
${}^{\prime}$: Statement 1: Journalist: the advice of social scientists is frequently overlooked by politicians making social policy. Statement 2: Because it is not unreasonable to discount scientific assertions, politicians should not generally be criticized for ignoring social science unless social scientists agree on the same claim, for social scientists, unlike physical scientists, seldom agree on the claims made even within their own specialty.
@@
Sufficient Assumption (6.6%)
Argument: Statement 1: Caffeine can kill or inhibit the growth of the larvae of several species of insects. Statement 2: One recent experiment showed that tobacco hornworm larvae die when they ingest a preparation that consists, in part, of finely powdered tea leaves, which contain caffeine. Statement 3: This result is evidence for the hypothesis that the presence of non-negligible quantities of caffeine in various parts of many diverse species of plants is not accidental but evolved as a defense for those plants.
@@
${}^{\prime}$: Statement 1: Caffeine produced for plant species’ own defense can kill or inhibit the growth of the larvae of several species of insects. Statement 2: One recent experiment showed that tobacco hornworm larvae die when they ingest a preparation that consists, in part, of finely powdered tea leaves, which contain caffeine. Statement 3: This result is evidence for the hypothesis that the presence of non-negligible quantities of caffeine in various parts of tobacco plant is not accidental but evolved as a defense for it.
@@
Strengthen (18.7%)
Argument: Statement1: In contemplating major purchases, businesses often consider only whether there is enough money left from monthly revenues after paying monthly expenses to cover the cost of the purchase. But many expenses do not occur monthly ; taking into account only monthly expenses can cause a business to overexpand. Statement2: So the use of a cash-flow statement is critical for all businesses.
@@
${}^{\prime}$: Statement 1: In contemplating major purchases, businesses often consider only whether there is enough money left from monthly revenues after paying monthly expenses to cover the cost of the purchase. But there are many expenses every month ; taking into account these expenses incorrectly can cause a business to overexpand. Statement 2: So the use of a cash-flow statement is critical for all businesses.
@@
Weaken (36.2%)
Argument: Statement1: The country of baurisia has, until now, been self-sufficient in both grain and meat. However, with growing prosperity in baurisia has come a steadily increasing per capita consumption of meat, and it takes several pounds of grain to produce one pound of meat. Statement2: Therefore, since per capita income in baurisia is almost certain to rise further but increases in domestic grain production are highly unlikely, baurisia is soon likely to become an importer of grain.
@@
${}^{\prime}$: Statement 1: The country of baurisia has, until now, been self-sufficient in both grain and meat. However, with growing prosperity in baurisia has come a steadily increasing per capita consumption of meat. Statement 2: Therefore, since per capita income in baurisia is almost certain to rise further but increases in domestic meat production are highly unlikely, baurisia is soon likely to become an importer of meat.
@@
Table 1: Example questions from the CLoMo benchmark, with the proportion of each logical relation.
@@ 2 Related Works
Complex reasoning has been highly concerned as a significant yet challenging task for inspecting advanced artificial intelligence. For example, for solving commonsense reasoning problems Talmor et al. (2019, 2021); Huang et al. (2019); Bhagavatula et al. (2020); Sap et al. (2019), the models Yasunaga et al. (2021, 2022); Liu et al. (2021); Huang et al. (2021b) are required to reasonable applying commonsense knowledge to conduct the reasoning process for the final answer. Furthermore, multi-step reasoning needs the models to perform multiple reasoning steps while maintaining consistency and faithfulness. To achieve this, synthetic compositional reasoning tasks Betz (2020); Tafjord et al. (2021); Han et al. (2022); Saparov and He (2023) incorporate first-order logic to inspect and improve models logical consistency Pan et al. (2023); Olausson et al. (2023); Sanyal et al. (2022); Yang and Deng (2021). Moreover, real-scenario compositional reasoning Yu et al. (2020); Liu et al. (2020); Dalvi et al. (2021); Huang et al. (2022) joins commonsense, consider the uncertainty of events in multi-step logical reasoning, which challenge current models Bao et al. (2023); Xu et al. (2023); Jiao et al. (2023, 2022); Huang et al. (2023a, 2021a) to solve real-world reasoning problems with faithfulness. Additionally, the more faithful models should be able to consider counterfactuals. For example, answering questions given counterfactual conditions Yu et al. (2023); Tandon et al. (2019); Qin et al. (2021), or narrating a counterfactual scenario Qin et al. (2019). However, previous studies on counterfactual reasoning barely pay attention to the logical consistency or faithfulness of models. Therefore, in this work, we propose the task of Counterfactual Logical Modification that challenges the model to satisfy a given logical relation restriction while generating counterfactuals.
@@
Reasoning poses a great challenge to LLMs. Currently, there is an increasing interest in the reasoning ability of LLMs. Evaluations are carried out from several perspectives, including mathematical reasoning, common-sense reasoning, logical reasoning, and domain-specific reasoning Chang et al. (2023); Zhong et al. (2023); Bang et al. (2023); Liu et al. (2023). However, most current methods of reasoning evaluation focus primarily on the accuracy of the final answer and neglect a comprehensive assessment of the reasoning process of LLMs. Such evaluation is not ideal for understanding the reasoning ability of models, as it ignores situations where models may obtain correct answers through unfaithful or spurious reasoning shortcuts Saparov and He (2023). Some recent research has started to evaluate the validity of the intermediate reasoning steps of LLMs Golovneva et al. (2022); Prasad et al. (2023). However, they mainly focus on the relationship between the intermediate step and the final answer rather than measuring whether the model understands the intermediate reasoning process. This paper proposes a novel logical reasoning benchmark that requires intermediate counterfactual thinking under logical relation restrictions. This leads to a more in-depth study of the model’s intermediate reasoning process.
@@ 3.1 Task Definition
A demonstration of the proposed Counterfactual Logical Modification is shown in Figure 1. An LLM is given the instruction as shown on the left-hand side, which can be illustrated by the diagram on the right-hand side. In the given instruction, Argument and Premise1 are related by a logical relation. We consider four main relations in practice, which are (R1) the premise provides a necessary assumption to the argument, (R2) the premise provides a sufficient assumption to the argument, (R3) the premise strengthen the argument, and (R4) the premise weaken the argument. The Argument and Premise1then constitute State 1 of the logical relation R. The additional Premise2 perturbs the logical relation R. The goal for the LLM is to maintain a State 2 with the given Premise2 and a modified ${}^{\prime}$ that R stands. To this end, it should properly edit the statements in Argument until the goal is reached.
@@ 3.2 Benchmark Construction
Since applicable data for the proposed task is lacking, we build a benchmark dataset by carefully selecting argument texts and collecting human annotation of the modified ${}^{\prime}$. We choose to use ReClor Yu et al. (2020) as the source data, considering that ReClor includes standardized multiple-choice questions on argument texts and logical relations from LSAT. We then recruit domain experts to conduct the annotation.
Given a ReClor question with context (the argument text), question, options, and the correct answer option, an annotator is required to provide a chosen wrong option (as the Premise2) and a corresponding modified context (i.e., the modified ${}^{\prime}$) to form a data point. The annotator is first instructed to read the whole question and comprehend the in-line logical relations, then choose one of the wrong options. After that, he/she edits the context by deleting, adding, or replacing text spans in the statements. The number of editions and the length of edited text spans are unrestricted as long as the statement partition is maintained.
We then post-process the question and the annotation so that for each data point, Argument comes from the original context, Premise1 comes from the correct option, and Premise2 and ${}^{\prime}$ come from the annotation.
4 SES: Self-Evaluation Scores
@@
Table 2: Prompts for LLM decomposed evaluation task.
@@
Human Performance 0.580 0.456 0.500 0.771 0.639
Table 3: Self-evaluation scores of LLM-generated counterfactual statements. NA: Necessary Assumption. SA: Sufficient Assumption. S: Strengthen. W: Weaken.
Counterfactual Modification Concept Graph Figure 2 To make the logic of counterfactual reasoning hold, we have a pair of primitive states $p_{1}$ (Premise1) and $a$ (Argument) from $p_{1}$ and $a$ satisfying the relation $r$, i.e., $\text{Pr}(r|p_{1},a)$ approaches $1$. The other claim $p_{2}$ (Premise2) and the modified ($a^{\prime}$) ${}^{\prime}$ satisfy the same relation $r$, i.e., $\text{Pr}(r|p_{2},a^{\prime})$ approaches $1$. In contrast, the relation between $p_{2}$ and the Argument should not satisfy the relation $r$, i.e., $\text{Pr}(r|p_{2},a)$ approaches 00.
Decomposed Self-Evaluation Score As the demonstrated complex logical reasoning ability of current large language models, we let the large language model estimate $\text{Pr}(r|p_{1},a)$, $\text{Pr}(r|p_{2},a^{\prime})$, and $\text{Pr}(r|p_{2},a)$, respectively. Specifically, we design a few binary classification tasks for the large language models, so that the probabilities are simplified to $c(r|p_{1},a)$, $c(r|p_{2},a^{\prime})$, and $c(r|p_{2},a)$$\in\{0,1\}$. Finally, the overall logical modification score is computed by:

### v4

@@ Abstract
In this study, we delve into the realm of counterfactual reasoning capabilities of large language models (LLMs). Our primary objective is to cultivate the counterfactual thought processes within LLMs and rigorously assess these processes for their validity. Specifically, we introduce a novel task, Counterfactual Logical Modification (CLoMo), and a high-quality human-annotated benchmark. In this task, LLMs must adeptly alter a given argumentative text to uphold a predetermined logical relationship. To effectively evaluate a generation model’s counterfactual capabilities, we propose an innovative evaluation metric, the decomposed Self-Evaluation Score (SES) to directly evaluate the natural language output of LLMs instead of modeling the task as a multiple-choice problem. Analysis shows that the proposed automatic metric aligns well with human preference. Our experimental results show that while LLMs demonstrate a notable capacity for logical counterfactual thinking, there remains a discernible gap between their current abilities and human performance. Code and data are available at https://github.com/Eleanor-H/CLOMO.
@@ 1 Introduction
In our study, we delve into the realm of evaluating large language models’ (LLMs) ability to generate counterfactually coherent thoughts. Figure 1 demonstrates the paradigm. Specifically, we proposed an innovative evaluation system that quantitatively measures the evolution of information in statement pairs, ensuring that they adhere to a specified logical relationship. Our approach includes designing a specialized task where models are presented with mismatched argument-premise pairs bound by a specific logical relation. The objective for these models is to adeptly modify the argument text until the specified logical relation is satisfactorily established. In conjunction with this task, we have created the first dataset of its kind, comprising dual argument-premise pairs, each annotated with a defined logical relation. This dataset is vital for facilitating logically restricted counterfactual modifications, and we have enriched it with human-written modifications to serve as a benchmark for evaluation.
Our experimental investigations encompass a range of large language models, including the latest GPT-4o [^1] https://openai.com/index/hello-gpt-4o/, GPT-4 OpenAI (2023), and GPT-3.5-Turbo OpenAI (2022), as well as smaller models from the LLaMA Touvron et al. (2023a) and LLaMA 2 Touvron et al. (2023b) families. Through these experiments, we have discerned that the task of CLoMo poses a significant challenge. It becomes evident that these models’ current counterfactual logical reasoning capabilities fall short of the desired proficiency. This observation underscores the need for further advancements in enhancing the counterfactual reasoning abilities of existing language models, paving the way for more sophisticated and logically coherent AI systems.
@@
  • We propose the task of Counterfactual Logical Modification and contribute a corresponding CLoMoto evaluate the counterfactual reasoning capability of LLMs in the scenario of complicated textual logical reasoning.
@@
  • We conduct experiments on LLMs (GPT-3.5, GPT-4) and small language models (the LLaMA and LLaMA 2 families) and find that CLoMo is a very challenging task and the counterfactual logical reasoning ability of the existing model needs to be improved.
@@
Necessary Assumption (38.5%) Argument: Statement1: Journalist: the advice of social scientists is frequently overlooked by politicians making social policy. Statement 2: Because it is not unreasonable to discount scientific assertions backed by weak evidence, politicians should not generally be criticized for ignoring social science, for social scientists, unlike physical scientists, seldom agree on the claims made even within their own specialty.
@@
Argument: Statement 1: Journalist: the advice of social scientists is frequently overlooked by politicians making social policy. Statement 2: Because it is not unreasonable to discount scientific assertions, politicians should not generally be criticized for ignoring social science unless social scientists agree on the same claim, for social scientists, unlike physical scientists, seldom agree on the claims made even within their own specialty.
@@
Sufficient Assumption (6.6%) Argument: Statement 1: Caffeine can kill or inhibit the growth of the larvae of several species of insects. Statement 2: One recent experiment showed that tobacco hornworm larvae die when they ingest a preparation that consists, in part, of finely powdered tea leaves, which contain caffeine. Statement 3: This result is evidence for the hypothesis that the presence of non-negligible quantities of caffeine in various parts of many diverse species of plants is not accidental but evolved as a defense for those plants.
@@
Argument: Statement 1: Caffeine produced for plant species’ own defense can kill or inhibit the growth of the larvae of several species of insects. Statement 2: One recent experiment showed that tobacco hornworm larvae die when they ingest a preparation that consists, in part, of finely powdered tea leaves, which contain caffeine. Statement 3: This result is evidence for the hypothesis that the presence of non-negligible quantities of caffeine in various parts of tobacco plant is not accidental but evolved as a defense for it.
@@
Strengthen (18.7%) Argument: Statement1: In contemplating major purchases, businesses often consider only whether there is enough money left from monthly revenues after paying monthly expenses to cover the cost of the purchase. But many expenses do not occur monthly ; taking into account only monthly expenses can cause a business to overexpand. Statement2: So the use of a cash-flow statement is critical for all businesses.
@@
Argument: Statement 1: In contemplating major purchases, businesses often consider only whether there is enough money left from monthly revenues after paying monthly expenses to cover the cost of the purchase. But there are many expenses every month ; taking into account these expenses incorrectly can cause a business to overexpand. Statement 2: So the use of a cash-flow statement is critical for all businesses.
@@
Weaken (36.2%) Argument: Statement1: The country of baurisia has, until now, been self-sufficient in both grain and meat. However, with growing prosperity in baurisia has come a steadily increasing per capita consumption of meat, and it takes several pounds of grain to produce one pound of meat. Statement2: Therefore, since per capita income in baurisia is almost certain to rise further but increases in domestic grain production are highly unlikely, baurisia is soon likely to become an importer of grain.
@@
Argument: Statement 1: The country of baurisia has, until now, been self-sufficient in both grain and meat. However, with growing prosperity in baurisia has come a steadily increasing per capita consumption of meat. Statement 2: Therefore, since per capita income in baurisia is almost certain to rise further but increases in domestic meat production are highly unlikely, baurisia is soon likely to become an importer of meat.
@@
Table 1: Example questions from the CLoMo benchmark, with the proportion of each logical relation. Counterfactual logical modifications regarding the change of state by a premise are highlighted.
@@ 2 Related Works
Complex reasoning has been highly concerned as a significant yet challenging task for inspecting advanced artificial intelligence. For example, for solving commonsense reasoning problems Talmor et al. (2019, 2021); Huang et al. (2019); Bhagavatula et al. (2020); Sap et al. (2019), the models Yasunaga et al. (2021, 2022); Liu et al. (2021); Huang et al. (2021b) are required to reasonable applying commonsense knowledge to conduct the reasoning process for the final answer. Furthermore, multi-step reasoning needs the models to perform multiple reasoning steps while maintaining consistency and faithfulness. To achieve this, synthetic compositional reasoning tasks Betz (2020); Tafjord et al. (2021); Han et al. (2022); Saparov and He (2023); Huang et al. (2024) incorporate first-order logic to inspect and improve models logical consistency Pan et al. (2023); Olausson et al. (2023); Sanyal et al. (2022); Yang and Deng (2021). Moreover, real-scenario compositional reasoning Yu et al. (2020); Liu et al. (2020); Dalvi et al. (2021); Huang et al. (2022) joins commonsense, consider the uncertainty of events in multi-step logical reasoning, which challenge current models Bao et al. (2023); Xu et al. (2023); Jiao et al. (2023, 2022); Huang et al. (2023a, 2021a) to solve real-world reasoning problems with faithfulness. Additionally, the more faithful models should be able to consider counterfactuals. For example, answering questions given counterfactual conditions Yu et al. (2023); Tandon et al. (2019); Qin et al. (2021), or narrating a counterfactual scenario Qin et al. (2019). However, previous studies on counterfactual reasoning barely pay attention to the logical consistency or faithfulness of models. Therefore, in this work, we propose Counterfactual Logical Modification that challenges the model to satisfy a given logical relation restriction while generating counterfactuals.
@@
Currently, there is an increasing interest in the reasoning ability of LLMs. Evaluations include several perspectives, such as mathematical reasoning, commonsense reasoning, logical reasoning, and domain-specific reasoning Chang et al. (2023); Zhong et al. (2023); Bang et al. (2023); Liu et al. (2023). However, most current reasoning evaluations focus primarily on the accuracy of the final answer and neglect a comprehensive assessment of the reasoning process. Such evaluation is not ideal for understanding the reasoning ability of models, as it ignores situations where models may obtain correct answers through unfaithful or spurious reasoning shortcuts Saparov and He (2023). Some recent research has started to evaluate the validity of the intermediate reasoning steps of LLMs Golovneva et al. (2022); Prasad et al. (2023). However, they mainly focus on the relationship between the intermediate step and the final answer rather than measuring whether the model understands the intermediate reasoning process. This paper proposes a novel logical reasoning benchmark that requires intermediate counterfactual thinking under logical relation restrictions. This leads to a more in-depth study of the model’s intermediate reasoning process.
@@ 3.1 Task Definition
A demonstration of the proposed Counterfactual Logical Modification is shown in Figure 1. An LLM is given the instruction as shown on the left-hand side, which can be illustrated by the diagram on the right-hand side. In the given instruction, Argument and Premise1 are related by a logical relation. We consider four main relations in practice, which are (R1) the premise provides a necessary assumption to the argument, (R2) the premise provides a sufficient assumption to the argument, (R3) the premise strengthen the argument, and (R4) the premise weaken the argument. The Argument and Premise1then constitute State 1 of the logical relation R. The additional Premise2 perturbs the logical relation R. The goal for the LLM is to maintain a State 2 with the given Premise2 and a modified Argument that R stands. To this end, it should properly edit the statements in Argument until the goal is reached. Table 1 lists R types, proportions, and sample questions.
@@ 3.2 Benchmark Construction
Given a data point [^2] Data source in Appendix C. with context (the argument text), question, options, and the correct answer option, an annotator is required to provide a chosen wrong option (as the Premise2) and a corresponding modified context (i.e., the modified Argument) to form a data point. The annotator is first instructed to read the whole question and comprehend the in-line logical relations, then choose one of the wrong options. After that, he/she edits the context by deleting, adding, or replacing text spans in the statements. The number of editions and the length of edited text spans are unrestricted as long as the statement partition is maintained.
We then post-process the question and the annotation so that for each data point, Argument comes from the original context, Premise1 comes from the correct option, and Premise2 and Argument come from the annotation.
[…]
@@
3.3 Data Statistics
Tables 2 and 3 demonstrate the dataset size, edit distance, and lengths of inputs of the gold inputs of CLoMo. CLoMo contains 1,000 manually constructed high-quality data points. According to Table 3, the input prompts of the chain-of-thought setting (CoT) have a medium of 379 tokens in the training set. The zero-shot setting (Zero) has a medium of 368 tokens in prompts, while the medium token length in the few-shot setting (Few) is up to 1,328 in the training set. Additionally, the output sequences Argument’ are a modification of Argument. The edit distance statistics in Table 2 shows that the most challenging data point in CLoMo has an edit distance of 66. The overall medium edit distance is 10, and the test set medium edit distance is 13. Therefore, CLoMo is a very challenging task.
Dataset Size Edit Distance
All Necessary Assumption Sufficient Assumption Strength Weaken Min Max Medium
Overall 1,000 385 66 187 362 1 66 10
Train 600 227 37 118 218 1 65 8
Dev 200 79 15 34 72 1 60 13
Test 200 79 14 35 72 1 66 13
Table 2: Statistics of CLoMo.
Max Min Mean Medium
CoT-Train 620 231 382.7 379
CoT-Dev 576 247 378.9 376
CoT-Test 620 216 374.7 371
Few-Train 1,569 1,180 1,331.7 1,328
Few-Dev 1,525 1,196 1,327.9 1,325
Few-Test 1,569 1,165 1,323.7 1,320
Zero-Train 609 220 371.7 368
Zero-Dev 565 236 367.9 365
Zero-Test 609 205 363.7 360
Table 3: CLoMo input length statistics by number of tokens. CoT: The chain-of-thought setting. Few: The few-shot setting. Zero: The zero-shot setting.
@@ 3.3 Data Statistics
Table 4: Prompts for LLM decomposed evaluation task in SES.
@@
GPT-4o1 0.680 0.696 0.667 0.800 0.625
Table 5: Self-evaluation scores (SES) of LLM-generated counterfactual statements. The backbone of SES is GPT-4. NA: Necessary Assumption. SA: Sufficient Assumption. S: Strengthen. W: Weaken.
4 SES: Self-Evaluation Scores
Counterfactual Modification Concept Graph Figure 2 demonstrates the graph. To make the logic of counterfactual reasoning hold, we have a pair of primitive states $p_{1}$ (Premise1) and $a$ (Argument) from $p_{1}$ and $a$ satisfying the relation $r$, i.e., $\Pr(r|p_{1},a)$ approaches $1$. The other claim $p_{2}$ (Premise2) and the modified $a^{\prime}$ (Argument) satisfy the same relation $r$, i.e., $\Pr(r|p_{2},a^{\prime})$ approaches $1$. In contrast, the relation between $p_{2}$ and the $a$ should not satisfy the relation $r$, i.e., $\Pr(r|p_{2},a)$ approaches 00.
Decomposed Self-Evaluation Score As the demonstrated complex logical reasoning ability of current large language models, we let the large language model estimate $\Pr(r|p_{1},a)$, $\Pr(r|p_{2},a^{\prime})$, and $\Pr(r|p_{2},a)$, respectively. Specifically, we design a few binary classification tasks for the large language models, so that the probabilities are simplified to $c(r|p_{1},a)$, $c(r|p_{2},a^{\prime})$, and $c(r|p_{2},a)$$\in\{0,1\}$. Table 4 demonstrates the prompts. The overall logical modification score is computed by: